You are considering a project with an initial cash outlay of ​$75,000 and expected cash flows of ​$21,750 at the end of each year for six years. The discount rate for this project is 9.7 percent. a.  What are the​ project's payback and discounted payback​ periods?

b.  What is the​ project's NPV?
c.  What is the​ project's PI?

d.  What is the​ project's IRR?
Answer #1

a. I have provided both cashflows and discounted cashflows in the below table

First find the discount factor like Year1=1/(1+9.7%), Year2=1/(1+9.7%)^2, Year3=1/(1+9,7%)^3 and so on.... Then multiply these discount factors with cashflows to find discounted cashflows

discount rate 9.70%
Cashflows Discount factor Discounted cashflows
Year0 -75000
Year1 21750 0.9116 19826.80
Year2 21750 0.8310 18073.66
Year3 21750 0.7575 16475.53
Year4 21750 0.6905 15018.71
Year5 21750 0.6295 13690.71
Year6 21750 0.5738 12480.14

Payback period is the time that takes to retain the initial cost of investment.

In 3 years, they receives 65,250 and remaining 9750 (75000-65250) in 4th year.

In 4th year=9750/21750=0.45

Payback period=3.45 years

For discounted Payback period, we should consider discounted cashflows.

In 4 years, they receives 69394.7 and remaining 5605.3 (75000-69394.7) in 5th year

in 5th Year=5605.3/13690.71=0.41

Discounted payback period=4.41 years

b. Use NPV function in EXCEL to calculate NPV

=NPV(rate,Year1 to Year6 cashflows)-Year0 cashflow

=NPV(9.7%,Year1 to Year6 cashflows)-75000

NPV=$20,565.56

To calculate PI, use same function

PI=NPV(rate,Year1 to Year6 cashflows)/Year0 cashflow

PI =NPV(9.7%,Year1 to Year6 cashflows)/75000

PI=1.27

Use IRR function

=IRR(Year0 to Year6 cashflows)

IRR=18.56%

discount rate 9.70%
Cashflows
Year0 -75000
Year1 21750
Year2 21750
Year3 21750
Year4 21750
Year5 21750
Year6 21750
NPV 20565.56
PI 1.27
IRR 18.56%